OpenStudy (anonymous):

Will someone please help me? Task 2—Non-Linear Systems of Equations Create a system of equations that includes one linear equation and one quadratic equation. Part 1. Show all work to solving your system of equations algebraically. Part 2. Graph your system of equations, and show the solution graphically to verify your solution.

OpenStudy (anonymous):

@phi

OpenStudy (phi):

How far did you get ?

OpenStudy (anonymous):

Not far, I don't understand it.

OpenStudy (anonymous):

f(x) = 2x + 1 y = x^2 + x - 2 I could use these two equations.

OpenStudy (phi):

Create a system of equations that means write down some equations (more than 1) one linear equation - that is the equation of a line and one quadratic equation. -- that is the equation of a parabola solve your system -- that means find the points where they line and parabola meet.

OpenStudy (anonymous):

Oh, that makes sense.

OpenStudy (phi):

I would make up equations that have nice answers. (and are easy to plot) first, does your quadratic "factor" ? i.e can you factor x^2 +x-2 ?

OpenStudy (phi):

yes, no or maybe?

OpenStudy (anonymous):

No?

OpenStudy (phi):

it factors \[ x^2 +x-2 =(x+2)(x-1) \] (You can find a nice parabola by starting with the factored form, and then using FOIL to multiply it out) the point is you have a reasonably easy parabola to plot. so we will use it. now let's find where the line meets the parabola. we do that by setting them equal f(x) = 2x + 1 y = x^2 + x - 2 2x+1 = x^2 + x - 2 add -2x - 1 to both sides

OpenStudy (phi):

2x+1 = x^2 + x - 2 add -2x - 1 to both sides the idea is to "move" all the terms to just one side.

OpenStudy (anonymous):

Okay I think I understand.

OpenStudy (phi):

what do you get after adding -2x-1 to both sides?

OpenStudy (anonymous):

-2x-1 = x?

OpenStudy (anonymous):

Hold on a second.

OpenStudy (phi):

2x+1 = x^2 + x - 2 put in -2x -1 on both sides 2x+1 -2x -1 = x^2 + x -2 -2x -1 can you simplify the left side? the left sides has 2x -2x and 1 - 1

OpenStudy (anonymous):

2(x + 1/2)?

OpenStudy (phi):

look just at 2x+1 -2x -1 you can swap the order, so it looks like this 2x - 2x + 1 - 1 what is 1-1 ?

OpenStudy (anonymous):

0

OpenStudy (phi):

notice when you see lots of terms, look for "just the numbers" and add them up (or subtract) now look at the terms with a single x in them 2x - 2x that is 2 x's take away 2 x's how many x's do we now have ?

OpenStudy (anonymous):

2

OpenStudy (phi):

2x - 2x that is 2 x's take away 2 x's

OpenStudy (anonymous):

Don't the lines connect at points (2,5) and (-1,-1)?

OpenStudy (phi):

not for this line. but we can tweak it. meanwhile, do you have the energy to finish this ?

OpenStudy (anonymous):

Yes, I need to get this done.

OpenStudy (phi):

ok, can you simplify 2x - 2x ?

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

or 0

OpenStudy (phi):

x^2 means x*x you have 2 x's take away 2 x's (it works the same as 2 spoons take away 2 spoons, or 2 horses take away 2 horses) you get 0 spoons, 0 horses, and 0 x's

OpenStudy (phi):

remember we set the equations equal \[ 2x+1 -2x -1 = x^2 + x -2 -2x -1 \\ 0 = x^2 + x -2 -2x -1 \] now try to simplify the right side. First do the pure numbers.